# MicroMoE: Fine-grained Load Balancing for Mixture-of-Experts with Token Scheduling

### v1

@@ 7.1. Experimental Setup
Baselines. We compare MicroMoE with four baseline systems.
@@
  • Megatron-LM (Megatron-LM) is a popular distributed training framework for large language models (LLMs). It supports various parallelism strategies, as well as state-of-the-art optimizations (recomputation).
@@
  • SmartMoE (SmartMoE) balances GPU loads by adjusting the expert placement within EP groups.
@@
  • FlexMoE (FlexMoE) achieves load balancing by dynamically adjusting replica counts based on expert loads. FlexMoE places expert replicas across the entire DP group, similar to MicroMoE’s asymmetric placement (§6.3).
@@
  • DeepSpeed (DeepSpeed) is a high-performance distributed framework for both LLM training and inference. We enable ZeRO-1 (ZeRO) optimization in DeepSpeed (currently, ZeRO-2 does not support PP, and Tutel (Tutel) does not support top-K¿1).
@@
We compare two variants of MicroMoE: “MicroMoE (w/o AR)” uses static, symmetric placement (§6.2), while “MicroMoE” uses adaptive, asymmetric placement (§6.3-6.4).
Figure 6. End-to-end speedup of different systems compared with Megatron-LM.
@@
Implementation. We implement MicroMoE upon Megatron-LM (Megatron-LM). MicroMoE provides a model wrapper similar to Pytorch’s Distributed Data Parallel (DDP) (DDP), enabling users to benefit from MicroEP’s fine-grained load balancing capabilities within their training jobs. We modify Megatron-LM, including its MoELayer and DDP with Python, and implement the token scheduling algorithm in MicroEP with C++.
@@
For fair comparison, we also implement SmartMoE and FlexMoE in Megatron-LM, as SmartMoE’s repository is outdated (last commit in 2023) (SmartMoE), and FlexMoE is not open-sourced (FlexMoE).
@@
Parallelization configurations. Due to the limited inter-node network bandwidth in our testbed, we only employ PP for inter-node parallelism. Specifically, we set the PP degree to the number of nodes used, and the DP degree to 8. We set the EP degree to 4, resulting in 2 EP groups per DP group. We set the parameter $d$ in MicroEP to 2, resulting in a single MicroEP group per DP group. We disable TP due to its high communication overhead.
@@
Other configurations. We use a small auxiliary loss (listed in the appendix) to prevent extreme load imbalance from degrading model accuracy. We enable the distributed optimizers in Megatron-LM, which resembles DeepSpeed’s ZeRO-1 (ZeRO). We disable the token dropping mechanism introduced by GShard (GShard). We use BF16 precision.
7.2. End-to-end Performance
@@
Figure 6 shows the end-to-end performance of all systems, varying models and number of GPUs. DeepSpeed exhibits poor performance with 16 or 32 experts. This is because DeepSpeed always adopts a padding mechanism, padding the load of each expert to the maximum expert load (GShard). This mechanism wastes significant time and memory when expert loads are highly imbalanced. With as few as 8 experts, the inefficiency of padding becomes less significant, allowing DeepSpeed to outperform Megatron-LM due to its system-level optimizations.
@@
Comparing SmartMoE and MicroMoE (w/o AR), where experts have uniform replica counts, MicroMoE (w/o AR) exhibits superior performance, thanks to MicroEP’s fine-grained token scheduling. While SmartMoE attempts to optimize performance by changing expert placement for load balancing among GPUs, it sometimes performs worse than vanilla Megatron-LM. This is because SmartMoE optimizes expert placement based on long-term expert load distributions. However, expert loads are highly dynamic during training, and SmartMoE’s long-term optimal placement may be sub-optimal for individual micro-batches.
@@
Comparing FlexMoE and MicroMoE (with AR), where experts have varied replica counts, MicroMoE exhibits superior performance, thanks to MicroEP. Compared with Megatron-LM, MicroMoE improves the end-to-end throughput by up to 47.6%, with an average improvement of 36.9%. The average performance improvement of MicroMoE surpasses FlexMoE, the second-best system, by 13.9%.
@@
Figure 7. Max GPU load normalized by average GPU load, varying skewness of expert loads. DP_degree=8, num_experts=32.
7.3. Load Balancing Capability
@@
We evaluate the load balancing capabilities of SmartMoE, FlexMoE, and MicroMoE with skewed expert loads. We generate expert loads following a Zipfian distribution with skewness $s$, where the probability of a token being assigned to the $i$-th most loaded expert is proportional to $i^{-s}$.
@@
Figure 7 shows the load balancing performance of different systems across varied skewness $s$. SmartMoE’s maximum GPU load increases as load skewness increases. While FlexMoE maintains relatively balanced GPU loads by adjusting expert replica counts, it falls short of achieving optimal load balance due to its lack of fine-grained dynamicity. MicroMoE (random) represents MicroMoE with pure random placement, which performs slightly worse than MicroMoE (w/o AR) with symmetric placement. MicroMoE (w/o AR) achieves perfect load balance when $s<1$ thanks to MicroEP’s fine-grained token scheduling. Nonetheless, its performance degrades when $s>1$ as uniform replica counts are insufficient for severe imbalances. MicroMoE with asymmetric placements can always achieve perfect load balance, due to the combination of both coarse-grained expert replacement and fine-grained token scheduling. Overall, MicroMoE exhibits the best load balancing capability among all systems and consistently achieves optimal load balance.
7.4. Execution Time Breakdown
@@
Figure 8. Execution time breakdown of an MoE layer. DP_degree=8, num_experts=32, micro_batch_size=8, sequence_length=2048, topK=2, hidden_size=4096, skewness ss=1.
@@
Figure 8 shows the execution time breakdown of an MoE layer across different systems. We omit DeepSpeed due to its poor performance. For all remaining systems, the primary bottleneck is expert computation time. MicroMoE achieves the shortest computation time by maintaining perfect load balance (with either symmetric or asymmetric placement, as shown in Figure 7).
@@
Specifically, the dispatch time consists of two primary components: (1) Preparation time, which includes the all-gather operation of expert load information and the scheduling of MicroEP. While MicroMoE introduces additional overhead in dispatch time due to token scheduling operations, we effectively minimize this impact through overlapping with other operations in Megatron-LM. (2) All-to-all communication time. Each all-to-all communication in dispatch and combine requires approximately 1.3 ms in Megatron-LM. We further evaluate dispatch time with DeepEP (DeepSeek-V3), a high-performance all-to-all backend. The results are presented in Appendix C.2.
7.5. Overhead Analysis
Scheduling Overhead.
@@ 7.5. Overhead Analysis
Figure 9. Scheduling time for MicroEP, varying number of experts and GPUs.
Figure 10. Migration time for adaptive replacement of MicroMoE.
@@
We evaluate the scheduling overhead of MicroEP, including the LPP solving time and token routing time. Our evaluation reveals that the LPP solving time is the dominant factor, which scales with the number of experts and GPUs. As shown in Figure 9, the scheduling overhead remains remarkably low, with a minimum time of approximately 100 us. Even with 64 GPUs and 256 experts, the scheduling time remains under 1 ms. This minimal overhead per micro-batch enables MicroEP to maintain high training throughput while providing load balancing benefits.
@@
Replacement overhead. The adaptive replacement strategy in MicroMoE necessitates model re-initialization to transition to new configurations. Although replacement is beneficial for load balancing, model re-initialization causes temporary suspension of training. Our evaluation focuses on two key components in the replacement overhead: the migration time of expert parameters and their optimizer states. As shown in Figure 10, the total migration time typically spans hundreds of milliseconds across different model configurations. These results emphasize the importance of carefully selecting the expert replacement frequency to optimize the trade-off between per-micro-batch training efficiency and overall replacement overhead.
7.6. Ablation Study
@@
Figure 11. Ablation study of dispatch time, with the same setting as Figure 8.
@@
We conducted an ablation study to evaluate three optimizations in MicroMoE: (1) warm solving in §5.1, (2) locality-aware routing in §5.2, and (3) overlapping in §5.4. All these optimizations aim to reduce the dispatch time. As shown in Figure 11, both warm solving and overlapping reduce scheduling time, while locality-aware routing reduces all-to-all communication time. With the combination of optimizations, MicroMoE introduces only 0.4 ms additional dispatch time compared to vanilla Megatron-LM. This modest overhead is substantially outweighed by the reduced computation time, ensuring the overall system efficiency of MicroMoE.
8. Related Work
@@
MoE training systems. Switch Transformer (Switch_Transformer) pioneers expert parallelism for transformer-based models. FairSeq (FairSeq) and FastMoE (FastMoE) provide PyTorch (PyTorch) plugins to optimize MoE training. DeepSpeed-MoE (DeepSpeed-MoE) combines EP with various optimizations, such as ZeRO (ZeRO) for large-scale MoE training. Tutel (Tutel) dynamically adapts parallelism strategies and pipeline degrees to handle dynamic workloads of MoE.
@@
MoE load balancing. Many studies propose algorithmic solutions to address the load imbalance issue of MoE. Shazeer et al. (shazeer2017outrageouslylargeneuralnetworks) introduce a load-balancing loss, which is further refined by subsequent works (Switch_Transformer; LocMoE; auxiliary-loss-free; attention_router; TA-MoE; DSelect-k). GShard (GShard) introduces the concept of expert capacity to limit the number of tokens assigned to each expert, which is adopted by later works (Switch_Transformer; LocMoE; Tutel). Some studies propose alternative gating algorithms in place of the original top-K gating strategy (BASE; expert_choice; Hash; LSH-MoE). As a systematic approach, MicroMoE is compatible with all these algorithmic solutions.
@@
Many studies offer systematic solutions for load balancing in MoE. FasterMoE (FasterMoE) models the time consumption of MoE and enables broadcasting the most popular expert to all devices. MoESys (MoESys) dynamically adjusts the number of training nodes to ensure load balance in multi-task scenarios. SmartMoE (SmartMoE) focuses on optimizing expert placement. FlexMoE (FlexMoE) adapts the number of replicas of each expert based on their popularity. Several other works (Hecate; SwiftMoE; Pro-Prophet; Prophet) embrace similar approaches to FlexMoE, and further reduce the adjustment overhead by integrating expert migration with ZeRO’s communication operations (ZeRO). While these solutions typically balance loads at a per-iteration, expert-level granularity, MicroMoE is capable of achieving optimal load balance at a per-micro-batch granularity.
@@
Janus (Janus) proposes a data-centric EP paradigm that avoids the inherent load imbalance issue in conventional expert-centric paradigms. However, Janus requires migrating all expert parameters within every micro-batch, which incurs substantial communication overhead, especially when dealing with large batch numbers and model scales.
9. Discussion
Inference. The load imbalance issue also exists in distributed inference systems with EP (DeepSpeed; PagedAttention; SGLang). Since inference jobs are more latency-sensitive than training, additional optimizations may be necessary to minimize the scheduling overhead of MicroMoE. For instance, we can replace the linear programming optimization with randomized algorithms or heuristic algorithms for reduced computational complexity, or predict the expert load distributions ahead of the gate network. We plan to employ MicroMoE for inference scenarios in future work.

### v2

@@ 7.1 Experimental Setup
Baselines. We compare FineMoE with four baseline systems.
@@
  • Megatron-LM [Megatron-LM] is a popular distributed training framework for large language models (LLMs). It supports various parallelism strategies, as well as state-of-the-art optimizations [recomputation].
@@
  • SmartMoE [SmartMoE] balances GPU loads by adjusting the expert placement within EP groups.
@@
  • FlexMoE [FlexMoE] achieves load balancing by dynamically adjusting replica counts based on expert loads. FlexMoE places expert replicas across the entire DP group, similar to FineMoE’s asymmetric placement (§6.3).
@@
  • DeepSpeed [DeepSpeed] is a high-performance distributed framework for both LLM training and inference. We enable ZeRO-1 [ZeRO] optimization in DeepSpeed (currently, ZeRO-2 does not support PP, and Tutel [Tutel] does not support top-K>1).
@@
We compare two variants of FineMoE: “FineMoE (w/o AR)” uses static, symmetric placement (§6.2), while “FineMoE” uses adaptive, asymmetric placement (§6.3-6.4).
@@
Implementation. We implement FineMoE upon Megatron-LM [Megatron-LM]. FineMoE provides a model wrapper similar to Pytorch’s Distributed Data Parallel (DDP) [DDP], enabling users to benefit from FineEP’s fine-grained load balancing capabilities within their training jobs. We modify Megatron-LM, including its MoELayer and DDP with Python, and implement the token scheduling algorithm in FineEP with C++.
@@
For fair comparison, we also implement SmartMoE and FlexMoE in Megatron-LM, as SmartMoE’s repository is outdated (last commit in 2023) [SmartMoE], and FlexMoE is not open-sourced [FlexMoE].
@@
Parallelization configurations. Due to the limited inter-node network bandwidth in our testbed, we only employ PP for inter-node parallelism. Specifically, we set the PP degree to the number of nodes used, and the DP degree to 8. We set the EP degree to 4, resulting in 2 EP groups per DP group. We set the parameter $d$ in FineEP to 2, resulting in a single FineEP group per DP group. We disable TP due to its high communication overhead.
@@
Other configurations. We use a small auxiliary loss (listed in the appendix) to prevent extreme load imbalance from degrading model accuracy. We enable the distributed optimizers in Megatron-LM, which resembles DeepSpeed’s ZeRO-1 [ZeRO]. We disable the token dropping mechanism introduced by GShard [GShard]. We use BF16 precision.
7.2 End-to-end Performance
@@
Figure 7 shows the end-to-end performance of all systems, varying models and number of GPUs. DeepSpeed exhibits poor performance with 16 or 32 experts. This is because DeepSpeed always adopts a padding mechanism, padding the load of each expert to the maximum expert load [GShard]. This mechanism wastes significant time and memory when expert loads are highly imbalanced. With as few as 8 experts, the inefficiency of padding becomes less significant, allowing DeepSpeed to outperform Megatron-LM due to its system-level optimizations.
@@
Comparing SmartMoE and FineMoE (w/o AR), where experts have uniform replica counts, FineMoE (w/o AR) exhibits superior performance, thanks to FineEP’s fine-grained token scheduling. While SmartMoE attempts to optimize performance by changing expert placement for load balancing among GPUs, it sometimes performs worse than vanilla Megatron-LM. This is because SmartMoE optimizes expert placement based on long-term expert load distributions. However, expert loads are highly dynamic during training, and SmartMoE’s long-term optimal placement may be sub-optimal for individual micro-batches.
@@
Comparing FlexMoE and FineMoE (with AR), where experts have varied replica counts, FineMoE exhibits superior performance, thanks to its fine-grained, per-micro-batch token scheduling.
In conclusion, compared with Megatron-LM, FineMoE improves the end-to-end throughput by up to 47.6%, with an average improvement of 36.9%. The average performance improvement of FineMoE surpasses FlexMoE, the second-best system, by 13.9%. Note that this is already the upper bound of performance improvement attainable through load balancing, since MicroMoE already achieves complete balance (detailed in § 7.3).
@@
Figure 8: Max GPU load normalized by average GPU load, varying skewness of expert loads. DP_degree=8, num_experts=32.
7.3 Load Balancing Capability
@@
We evaluate the load balancing capabilities of SmartMoE, FlexMoE, and FineMoE with skewed expert loads. We generate expert loads following a Zipfian distribution with skewness $s$, where the probability of a token being assigned to the $i$-th most loaded expert is proportional to $i^{-s}$.
@@
Figure 8 shows the load balancing performance of different systems across varied skewness $s$. SmartMoE’s maximum GPU load increases as load skewness increases. While FlexMoE maintains relatively balanced GPU loads by adjusting expert replica counts, it falls short of achieving optimal load balance due to its lack of fine-grained dynamicity. FineMoE (random) represents FineMoE with pure random placement, which performs slightly worse than FineMoE (w/o AR) with symmetric placement. FineMoE (w/o AR) achieves perfect load balance when $s<1$ thanks to FineEP’s fine-grained token scheduling. Nonetheless, its performance degrades when $s>1$ as uniform replica counts are insufficient for severe imbalances. FineMoE with asymmetric placements can always achieve complete load balance, due to the combination of both coarse-grained expert replacement and fine-grained token scheduling. Overall, FineMoE exhibits the best load balancing capability among all systems and consistently achieves complete load balance.
7.4 Execution Time Breakdown
@@
Figure 9: Execution time breakdown of an MoE layer. DP_degree=8, num_experts=32, micro_batch_size=8, sequence_length=2048, topK=2, hidden_size=4096, skewness ss=1.
@@
Figure 9 shows the execution time breakdown of an MoE layer across different systems. We omit DeepSpeed due to its poor performance. For all remaining systems, the primary bottleneck is expert computation time. FineMoE achieves the shortest computation time by maintaining perfect load balance (with either symmetric or asymmetric placement, as shown in Figure 8).
@@
Specifically, the dispatch time consists of two primary components: (1) Preparation time, which includes the all-gather operation of expert load information and the scheduling of FineEP. While FineMoE introduces additional overhead in dispatch time due to token scheduling operations, we effectively minimize this impact through overlapping with other operations in Megatron-LM. (2) All-to-all communication time. Each all-to-all communication in dispatch and combine requires approximately 1.3 ms in Megatron-LM.
Figure 10: Dispatch time of FineEP and EP with DeepEP and NCCL, varying number of GPUs.
@@
Figure 11: Scheduling time for FineEP, varying number of experts and GPUs.
Figure 12: Migration time for adaptive replacement of FineMoE.
7.5 Inter-node Communication with DeepEP
We evaluate the dispatch time of FineEP and vanilla EP for inter-node communication. We additionally integrate FineEP with DeepEP [DeepSeek-V3], a high-performance all-to-all communication backend. Currently, Megatron-LM [Megatron-LM] supports both NCCL [NCCL] (by default) and DeepEP for all-to-all communication.
Experimental considerations. Due to testbed limitations, two important experimental considerations should be noted: (1) Our testbed consists of 8 GPUs but only 2 NICs per node, resulting in limited inter-node network bandwidth. Therefore, we avoid employing EP or FineEP across multiple nodes in §7. However, since this section focuses on evaluating the performance of different communication backends, we expand the communication group to multiple nodes. Consequently, the all-to-all time for inter-node communication is significantly higher than the intra-node communication. (2) §7 focuses on system performance, so we compare FineEP using 8 GPUs per group with EP using 4 GPUs per group ($d=2$). However, this section focuses on communication performance, so we compare FineEP and EP using the same group size.
Overhead of inter-node communication. Notably, since FineEP expands the all-to-all group size by a factor of $d$, FineEP may convert some inter-node communication into inter-node, leading to extra overhead. However, this overhead is minimal in two typical scenarios: (1) When $d\times\text{EP\_degree}\leq$ # GPUs per node, the all-to-all communication in FineEP remains entirely intra-node. (2) When EP_degree is super large (e.g., 64 in DeepSeek-v3 [DeepSeek-V3]), nearly all communication is inherently inter-node. Consequently, FineEP incurs negligible overheads in both single-node and massive-node scenarios. Furthermore, our communication-aware scheduling mechanism can jointly optimize the time of both communication and computation, as shown in Appendix C.2.
Results. Figure 16 shows the dispatch time comparison between FineEP and EP using both DeepEP and NCCL, varying number of GPUs. We use the same setting as in §7.4, except for the all-to-all group size. DeepEP exhibits better performance than NCCL due to its superior implementation. When using NCCL, FineEP requires less time than EP, thanks to the locality-aware routing in §5.2. However, when using DeepEP, FineEP requires more time than EP due to data format incompatibilities between DeepEP and Megatron-LM. Consequently, Megatron-LM needs to pre-process the data for DeepEP, while FineEP incurs a higher pre-processing overhead than EP. We believe that FineMoE will yield lower communication overheads on other practical testbeds (e.g., with one NIC per GPU).
7.6 Overhead Analysis
@@
Scheduling Overhead. We evaluate the scheduling overhead of FineEP, including the LPP solving time and token routing time. Our evaluation reveals that the LPP solving time is the dominant factor, which scales with the number of experts and GPUs. As shown in Figure 12, the scheduling overhead remains remarkably low, with a minimum time of approximately 100 us. Even with 64 GPUs and 256 experts, the scheduling time remains below 1 ms. This minimal overhead per micro-batch enables FineEP to maintain high training throughput while providing load balancing benefits. Additionally, we evaluate the performance of pipelining to hide the scheduling latency, which results are shown in Appendix C.3.
@@
Replacement overhead. The adaptive replacement strategy in FineMoE necessitates model re-initialization to transition to new configurations. Although replacement is beneficial for load balancing, model re-initialization causes temporary suspension of training. Our evaluation highlights two key components in the replacement overhead: the migration time of expert parameters and their optimizer states. As shown in Figure 12, the total migration time typically spans hundreds of milliseconds across different model configurations.
The above results emphasize the importance of carefully selecting the expert replacement frequency to optimize the trade-off between per-micro-batch training efficiency and overall replacement overhead. In practice, we recommend tuning the replacement interval to 50 iterations during the beginning phase of training, which adds less than 1% overhead to the entire system. One may increase this interval to several hundred iterations or even make no replacement when workloads become less volatile, as shown in Figure 2.
@@
Figure 13: Ablation study of dispatch time, with the same setting as Figure 9.
7.7 Ablation Study
@@
We conducted an ablation study to evaluate three optimizations in FineMoE: (1) warm solving in §5.1, (2) locality-aware routing in §5.2, and (3) overlapping in §5.4. All these optimizations aim to reduce the dispatch time. As shown in Figure 13, both warm solving and overlapping reduce scheduling time, while locality-aware routing reduces all-to-all communication time. With the combination of optimizations, FineMoE introduces only 0.4 ms additional dispatch time compared to vanilla Megatron-LM. This modest overhead is substantially outweighed by the reduced computation time, ensuring the overall system efficiency of FineMoE.
8 Related Work
@@
MoE training systems. Switch Transformer [Switch_Transformer] pioneers expert parallelism for transformer-based models. FairSeq [FairSeq] and FastMoE [FastMoE] provide PyTorch [PyTorch] plugins to optimize MoE training. DeepSpeed-MoE [DeepSpeed-MoE] combines EP with various optimizations, such as ZeRO [ZeRO] for large-scale MoE training. Tutel [Tutel] dynamically adapts parallelism strategies and pipeline degrees to handle dynamic workloads of MoE.
@@
MoE load balancing. Many studies propose algorithmic solutions to address the load imbalance issue of MoE. Shazeer et al. [shazeer2017outrageouslylargeneuralnetworks] introduce a load-balancing loss, which is further refined by subsequent works [Switch_Transformer, LocMoE, auxiliary-loss-free, attention_router, TA-MoE, DSelect-k]. GShard [GShard] introduces the concept of expert capacity to limit the number of tokens assigned to each expert, which is adopted by later works [Switch_Transformer, LocMoE, Tutel]. Some studies propose alternative gating algorithms in place of the original top-K gating strategy [BASE, expert_choice, Hash, LSH-MoE]. As a systematic approach, FineMoE is compatible with all these algorithmic solutions.
@@
Many studies offer systematic solutions for load balancing in MoE. FasterMoE [FasterMoE] models the time consumption of MoE and enables broadcasting the most popular expert to all devices. MoESys [MoESys] dynamically adjusts the number of training nodes to ensure load balance in multi-task scenarios. SmartMoE [SmartMoE] focuses on optimizing expert placement. FlexMoE [FlexMoE] adapts the number of replicas of each expert based on their popularity. Several other works [Hecate, SwiftMoE, Pro-Prophet, Prophet] embrace similar approaches to FlexMoE, and further reduce the adjustment overhead by integrating expert migration with ZeRO’s communication operations [ZeRO]. While these solutions typically balance loads at a per-iteration, expert-level granularity, FineMoE is capable of achieving optimal load balance at a per-micro-batch granularity.
LPLB [LPLB] is a recent work that leverages linear-programming-based token scheduling for MoE load balancing, similar to MicroMoE. However, it lacks comprehensive analysis and experimental evaluation.  [^4] By the time we wrote this paper, LPLB was still in the early research stage and had no evaluation results. As far as we know, LPLB is the only relevant work that schedules tokens for load balancing. In contrast, this paper theoretically analyzes the relationship between expert placement and token scheduling, proposes multiple placement strategies, and incorporates key optimizations such as locality-aware routing. We believe that our work provides clearer insights and practical guidance for designing next-generation MoE systems.
@@
Janus [Janus] proposes a data-centric EP paradigm that avoids the inherent load imbalance issue in conventional expert-centric paradigms. However, Janus requires migrating all expert parameters within every micro-batch, which incurs substantial communication overhead, especially when dealing with large batch numbers and model scales.
9 Discussion